### v2

# Ultraviolet photon production rates of the first stars: Impact on the He II 1640 Å emission line from primordial star clusters and the 21-cm signal from cosmic dawn

Abstract
The first stars, the chemically pristine Population III, likely played an important role in heating the intergalactic medium during the epoch of cosmic dawn. The very high effective temperatures ( K) predicted for the most massive Population III stars could also give rise to tell-tale signatures in the emission-line spectra of early star clusters or small galaxies dominated by such stars. Important quantities in modelling their observational signatures include their photon production rates at ultraviolet energies at which photons are able to ionize hydrogen and helium, dissociate molecular hydrogen and cause Lyman- heating. Here, we model the spectral energy distributions of Population III stars to explore how these key quantities are affected by the initial mass and rotation of Population III stars given a wide range of models for the evolution of these stars. Our results indicate that rotating Population III stars that evolve to effective temperatures K could potentially give rise to a very strong HeII 1640 Å emission line in the spectra from primordial star clusters, without requiring stellar masses of indicated by previous models for non-rotating Population III stars. The observable impact on 21-cm signatures from cosmic dawn and the epoch of reionization from our set of rotating stars that evolve to K is modest, except in case of high Population III star formation efficiencies which imprint potentially detectable features in the global 21-cm signal and 21-cm power spectrum.
keywords:
Stars: Population III – stars: rotation – cosmology: dark ages, reionization, first stars – galaxies: high-redshift1 Introduction
The first generation of stars, the chemically pristine Population III (hereafter Pop III), is believed to have played an important role in heating the early intergalactic medium during the epoch of cosmic dawn, in producing the first elements beyond lithium, and in the production of early black holes (for a recent review, see Klessen and Glover, 2023). Such stars probably started forming around redshift , about 100 Myr after the Big Bang, but may have continued to form in pockets of primordial gas throughout the reionization epoch and well into the post-reionization era (e.g. Johnson, 2010; Mebane et al., 2018; Liu and Bromm, 2020). In terms of observational probes, 21-cm cosmology may be able to probe the nature of Pop III stars at –30 through their impact on the temperature state of the intergalactic medium (e.g. Pochinda et al., 2024; Gessey-Jones et al., 2025; Ventura et al., 2025). At somewhat lower redshifts (–15) one may hope to detect Pop III signatures in individual objects, either among samples of gravitationally lensed, high-redshift stars (Windhorst et al., 2018; Zackrisson et al., 2024), in gravitationally lensed Pop III star clusters and galaxies (Zackrisson et al., 2012; Vanzella et al., 2023; Fujimoto et al., 2025), or within galaxies where Pop III star formation is able to continue alongside metal-enriched star formation (e.g. Sarmento et al., 2018; Venditti et al., 2023; Maiolino et al., 2024; Wang et al., 2024). Other ways of empirically studying the properties of Pop III stars include the chemical signatures of Pop III supernovae in gas absorption systems (e.g. D’Odorico et al., 2023; Sodini et al., 2024; Vanni et al., 2024), in high-redshift quasars (Yoshii et al., 2022) and in second-generation stars in the local Universe (e.g. Fraser et al., 2017; Ishigaki et al., 2018; Jiang et al., 2024; Bonifacio et al., 2025), as well as the direct detection of supernovae (e.g. Moriya et al., 2022; Venditti et al., 2024), gamma-ray bursts (e.g. Lazar and Bromm, 2022), and tidal disruption events (Kar Chowdhury et al., 2024) produced by Pop III stars.
Due to their chemically pristine composition, Pop III stars are expected to be more massive (e.g. Klessen and Glover, 2023), and reach higher main-sequence temperatures (e.g. Schaerer, 2002) than their metal-enriched descendants. This gives rise to higher hydrogen- and helium-ionizing photon production rates compared to metal-enriched stars, and also differences in the Lyman–Werner (LW; -dissociating) and the Lyman-band rates important for the Wouthuysen–Field coupling during cosmic dawn. However, the production of photons in these different energy regimes depends not only on initial mass but also on stellar rotation (e.g. Yoon et al., 2012; Murphy et al., 2021b; Sibony et al., 2022; Liu et al., 2025a, b; Hawcroft et al., 2025), and simulations have indicated that at least some Pop III may rotate at close to break-up speed (Hirano and Bromm, 2018).
Here, we use the Muspelheim (Zackrisson et al., 2024) models for the spectral energy distributions (SEDs) of Pop III stars to explore the impact of different sets of stellar evolutionary tracks on the ultraviolet (UV) radiation output of Pop III stars. These models use stellar atmosphere SEDs instead of the blackbody SEDs adopted in many previous studies (e.g. Yoon et al., 2012; Murphy et al., 2021b; Klessen and Glover, 2023). In section 2, we present these models and in section 3 we report the resulting hydrogen- and helium-ionizing photon production rates, the LW and the Ly-band rates, along with a comparison to results derived from blackbody SEDs. The impact on studies of the UV SEDs of Population III star clusters and galaxies is explored in Section 4 and the impact on 21 cm-cosmology is shown in Section 5. Various future avenues for improvement of the models are discussed in section 6. Section 7 summarises our findings.
2 SED models of Population III stars
The SED models for Pop III stars (Muspelheim v.1) presented by Zackrisson et al. (2024) couple sets of stellar evolutionary tracks to a grid of stellar atmospheres to produce SEDs as a function of Zero Age Main Sequence (ZAMS) mass and age. Important simplifications used in the Muspelheim v.1 models include the use of stellar atmospheres for spherical stars without mass loss, and with fixed primordial abundances of hydrogen and helium. Hence, these models neglect the non-sphericity that come from high rotation rates, the effects of winds, and the effects of surface helium enhancement and surface self-pollution of metals during the course of evolution. These shortcomings are further discussed in Section 6. In this paper, we make use of Muspelheim SED models based on selected Pop III tracks from Yoon et al. (2012) at –500 M⊙, Windhorst et al. (2018) at –100 M⊙, Murphy et al. (2021a) at –120 M⊙, Volpato et al. (2023) at –1000 M⊙ and Costa et al. (2025) at –200 M⊙. Out of these sets, the Yoon et al. (2012) and Murphy et al. (2021a) sets feature models both with and without rotation. Although Yoon et al. (2012) present stellar evolutionary tracks for several different initial velocities in the range 0.0–0.6 (where is the Keplerian velocity at the equatorial surface), we here limit the discussion to the and 0.4 models, for simplicity.
While many of these tracks are similar close to the ZAMS, they differ significantly in their predictions of their low- extension at late stages of stellar evolution. While Zackrisson et al. (2024) argue that late evolution to K have important consequences for the prospects of detecting lensed, individual Pop III stars at high redshifts, one could expect that such differences would be minor when it comes to the UV photon production rates of Pop III stars, since these are dominated by the more long-lived high- stages of these stars. However, the Yoon et al. (2012) tracks for rotating stars, which also include the effects of magnetic fields, differ from many other models of rotating Pop III stars in the literature (e.g. Murphy et al., 2021a; Sibony et al., 2022; Martinet et al., 2023; Volpato et al., 2024; Tsiatsiou et al., 2024) in that some of the stars with high initial rotation velocity evolve to very high ( K) at bolometric luminosities an order of magnitude above their ZAMS luminosities at the end of their lifetimes (similar effects are, however, seen in the rotating Pop III models in the recent paper by Hassan et al., 2025), with significant impact on their UV properties.
In this paper, we explore the UV properties of Pop III stars in terms of their production of H-, He- and He+-ionizing photons, as well as their production of LW and Lyman-band photons.
The instantaneous photon production rates , expressed in units of photons per second, are defined as:
where is the monochromatic luminosity (erg-1 s-1 Å-1) at wavelength , is Planck’s constant and the speed of light. The integration limits for the different rates are: and Å for ; and Å for ; and Å for ; Å and Å for ; Å and Å for . We also present , the Lyman-band photon production count per baryon, integrated over the stellar lifetime , given by:
where is the mass of a proton.
3 Ionizing and Lyman–Werner fluxes
As previously shown by e.g. Bromm et al. (2001), Larkin et al. (2023), and Liu et al. (2025a), the SEDs of stellar atmosphere models predict UV properties of hot Pop III stars that deviate from those of blackbody SEDs scaled to the same effective temperatures and bolometric luminosities . We illustrate this in Fig. 1, where we compare the SEDs of the TLUSTY stellar atmosphere models used in Muspelheim to those of blackbody SEDs for stars at – K. The photon counts per wavelength () differ between blackbody and stellar atmosphere SEDs in all the UV bands, albeit to different extent and in different directions throughout the range of model SEDs shown. It is worth noting, that even though the highest- SED shown ( K, a temperature reached only by rotating Yoon et al. (2012) 20–150 M⊙ Pop III stars in our set) exhibits the most blackbody-like shape (with small continuum breaks only), the peak of the stellar-atmosphere SED occurs at shorter wavelengths compared to the corresponding blackbody, an effect previously seen in the stellar atmosphere SEDs for Pop III stars by Bromm et al. (2001) and also for pure hydrogen atmospheres for the very hot central stars of planetary nebulae (Rauch, 2003).
The effect on the ionizing- and LW photon production rates can be seen in Fig. 2 where we compare results generated from blackbody SEDs (dashed lines; in good agreement with the results presented by Murphy et al. 2021b) and stellar atmospheres (solid lines) for the non-rotating Murphy et al. (2021a) stellar evolutionary tracks with initial masses in the range 9–120 M⊙. The general behaviour follows what we would expect from Fig. 1, as the shape111But not the scaling, which in Fig. 1 has been chosen to avoid overlap of the SEDs in the plot of the K, and SEDs in Fig. 1 are representative of early evolution along the , and Murphy et al. (2021a) tracks in Fig. 2. The blackbody SEDs have consistently higher H-ionizing rates, but lower He-ionizing rates, for all the stellar masses plotted, although by no more than -0.3 dex. The behaviour of the He+-ionizing rates is more complex, as the blackbody under-predicts the flux at 85-120 M⊙ by -0.3 dex, but then severely starts to over-predict the He+-ionizing flux by up to dex at lower masses due to the presence of a strong continuum break in the stellar atmosphere SEDs at the He+ edge (seen in the K and K SEDs in Fig. 1). In the case of the LW flux, the blackbody SED over-predicts the flux by up to dex at the highest masses (as seen in the K SED), becomes nearly identical at 15 M⊙ (as seen in the K SED) and drops below the stellar atmosphere SED by dex by 9 M⊙ (as seen in the K SED).
When modelling the impact of Pop III stars on the intergalactic medium, the detailed temporal evolution of the UV photon rates is often neglected in favour of lifetime-integrated quantities (for a counterexample, see Gessey-Jones et al., 2022). This may be reasonable in cases where the lifetimes of the Pop III stars are short compared to timescales related to the thermal evolution of the IGM or changes in cosmic star formation rate density, as is the case for the more massive Pop III stars. In Fig. 3 we present the lifetime-integrated photon production rates of our Pop III models per solar mass, compared to blackbody models from Klessen and Glover (2023) and models for rotating stars with homogeneous chemical evolution from Liu et al. (2025a, further discussed in section 5). For stellar masses , all models except the Yoon et al. (2012) and Liu et al. (2025a) rotating models agree on the H-ionizing and He-ionizing rates to within dex. The Yoon et al. (2012) models at 20 MZAMS 200 M⊙, and all of the Liu et al. (2025a) models, undergo chemically homogeneous evolution which significantly boosts the ionizing photon production of these stars. At lower masses, models based on blackbodies are seen to deviate from those based on stellar atmosphere SEDs by up to dex for H and dex for He.
In the case of the He+-ionizing photon production rates, models based on blackbody SEDs at consistently produce higher values than corresponding models based on stellar atmosphere SEDs, as expected from Fig. 2, with an offset that reaches dex at the lowest masses. Models for rotating Yoon et al. (2012) stars at 20–300 M⊙ produce higher He+-ionizing rates than all other models at comparable masses. This increase of He+-ionzing photons due to rotation was already highlighted by Yoon et al. (2012), although the increase compared to non-rotating models becomes much higher in models based on stellar atmosphere SEDs (peaking at dex at 20 M⊙) compared to the blackbody SEDs ( dex) used by Yoon et al. (2012), as shown in their fig. 7. Liu et al. (2025a) is not included in the He+ panel as their spectra are cut of at 124Å which leaves out a significant portion of the He+-ionizing photons, especially at high Teff (Fig. 1).
For the LW band, models that undergo CHE produce somewhat lower rates (–0.3 dex) than other models at 50 M⊙ while lower masses produce rates on the higher side compared with other models. The Klessen and Glover (2023) blackbody models also produce systematically lower rates at low stellar masses (reaching offsets of dex at 1.7 M⊙) compared to models based on stellar atmosphere SEDs. At high masses, the most notable outliers are the models by Schaerer (2002), which at produce much higher rates than all other models (up to dex higher). This discrepancy is discussed further in Section 6.2.
The lifetime-average photon production rates corresponding to the lifetime-integrated rates in Fig. 3 for our Yoon et al. (2012) and Murphy et al. (2021a) models can be found in Table 1. tables and both lifetime-integrated and lifetime-average photon production rates for all models can be downloaded from the Muspelheim webpage (see Section Data Availability).
4 Impact on the the HeII1640 emission line from Population III star clusters
The very high He+-ionizing photon production rates of Pop III rotating Yoon et al. (2012) stars are interesting in the context of HeII emission lines from high-redshift galaxies. In particular, the HeII1640 emission line has been proposed as a potential Pop III signature, as it can become very strong (with rest-frame equivalent widths up to Å) in the presence of K stars (e.g. Tumlinson and Shull, 2000; Schaerer, 2002; Raiter et al., 2010; Nakajima and Maiolino, 2022). For non-rotating stars, effective temperatures required to produce Å are not reached until , which means that a very strong HeII1640 line from an integrated stellar population (Pop III star cluster or Pop III-dominated galaxy) could serve as an indicator of a top-heavy Pop III stellar initial mass function (IMF). Recently, a system with Å and no detectable metal lines, located in the halo of the galaxy GNz-11 at , was found by Maiolino et al. (2024) who interpreted this as evidence of a Pop III IMF extending at least up to 500 M⊙ and a total mass of in Pop III stars.
In Fig. 4, we explore the temporal evolution of the He+-ionizing rates of our Yoon et al. (2012) models, with and without rotation. Since rotation in these models shifts the early main sequence to lower and slightly lower bolometric luminosities, rotation leads to lower He+ rates during early evolution (although this phase is very brief in the case of the 500 M⊙ model). However, since the rotating stars eventually evolve to become hotter and more luminous than their non-rotating counterparts for , the He+-ionizing photon rates of these rotating stars grow by several orders of magnitude at high ages. This effect, coupled to the extended lifetimes of these stars (due to rotational mixing renewing the fuel in the core), explains the boosted time-integrated rates for rotating Yoon et al. (2012) models shown in Fig. 3.
Since Fig. 4 indicates that the He+-ionizing rates of these 20–30 M⊙ rotating models can become comparable, for about Myr at the end of their lifetimes, to the peak He+-ionizing rates of 100-150 M⊙ stars (maintained over a similar timescale, but during early evolution), it seems that very strong HeII1640 emission in principle could be produced by rotating Pop III stars without the presence of stars.
To explore this, we use the latest version of the photoionization code cloudy (v23.01), last described by Chatzikos and others (2023) and Gunasekera and others (2023), to investigate the effect of gas parameters on the line strength and equivalent width of the HeII1640 emission line in gas clouds surrounding single stars from the Yoon et al. (2012) set. We assume primordial gas abundances, a plane-parallel nebula, and an ionization parameter in the range to .
In Fig. 5 we show the resulting equivalent widths of the HeII1640 emission line for a subset of Yoon et al. (2012) tracks. Comparing our results with HeII1640 EWs from Schaerer (2003) and Raiter et al. (2010) – which both assume fully sampled Pop III IMFs – we are able to obtain only slightly higher HeII1640 equivalent widths than the most extreme values seen in their models. However, the major difference is that both Schaerer (2003) and Raiter et al. (2010) require IMFs dominated by very massive stars to reach the high EWs that we are able to obtain for 20 M⊙ stars. We also explore other ionization parameters, not shown in the figure, and find that the equivalent widths are largely stable at log(U) -3, but note a significant drop at log(U) = -4 in agreement with Raiter et al. (2010).
Our results indicate that a stellar population of Yoon et al. (2012) rotating stars with an IMF peaking at in principle could produce equally strong HeII1640 equivalent widths as what has been seen for Pop III models with non-rotating stars and more top-heavy IMFs (like the 50-500 , Salpeter-slope model, and the , , 1–500 M⊙ lognormal model from Raiter et al. 2010). Invoking rotating Yoon et al. (2012) Pop III stars to explain the HeII1640 emission in the halo of GNz-11 would not only remove the requirement of Pop III stars to explain the high equivalent width, but also slightly reduce the total mass in Pop III stars required to explain the line luminosity. Because our rotating stars reach somewhat higher ratio than the models used by Maiolino et al. (2024), we find that the total mass in Pop III stars could be reduced by a factor of compared to their estimate, although this would happen only in a scenario where all Pop III stars within the system are rapidly rotating stars caught right in their hottest phase. An IMF strongly peaked around is not favoured by current simulations (for a review, see Klessen and Glover, 2023), but the chemical properties of several extremely metal-poor stars in the Milky Way halo and in ultrafaint dwarfs can be explained by supernovae in the mass range, either from faint supernovae with fallback, hypernovae or a combination of supernovae and winds from rapidly rotating Pop III stars (e.g. Ishigaki et al., 2018; Skúladóttir et al., 2021; Placco et al., 2021; Jeena et al., 2023). Under the assumption of second-generation stars having been enriched by single Pop III supernovae, the exercise of fitting models of supernova ejecta to stellar abundance patterns can give the appearance of a log-normal Pop III IMF peaked around Ishigaki et al. (2018), but folding in duo-enrichment and the propensity for failed supernovae can also make observed abundance patterns consistent with an IMF extending to much lower and higher masses (Jiang et al., 2024). Finally, a Pop III IMF peaked around 20 is also consistent with the Pop III IMF range found by cosmological models of Milky Way formation, which self-consistently follow the chemical evolution across cosmic time and constrain the Pop III IMF by matching the number and properties of observed ancient metal-poor stars (e.g. Koutsouridou et al., 2024).
It is important to note that since the K phase of the 20 M⊙ rotating Yoon et al. (2012) model is reached only at the very end of its Myr lifetime, a Pop III system that also hosts very massive Pop III stars with -–260 (and lifetimes as short as Myr for the models considered here) will already have experienced the explosion of pair-instability supernovae (e.g. Heger and Woosley, 2002). If the hosting halo has low mass, these energetic explosions might have cleared away much of the remaining gas from the vicinity of the stars, which would lead to very low ionization parameters and thus reduced HeII1640 equivalent widths. For the lowest-mass Pop III haloes, a large fraction of the gas may even be removed from the halo (e.g. Bromm et al., 2003; Kitayama and Yoshida, 2005; Mead et al., 2025). This would preclude the detection of any emission lines associated with the late evolution of these 20 M⊙ stars, unless the gas can quickly recollapse (in the Magg et al. 2022b simulations of Pop III haloes, recollapse can happen in as little as Myr, although the median recollapse time is 50 Myr). Even without supernovae, the initially dense gas around a cluster of Pop III stars in haloes may be expelled to large radii and low densities in response to the radiation field from the stars, with long recombination time-scales and significant leakage of ionizing photons into the intergalactic medium as a result (e.g. Sibony et al., 2022). Detailed simulations would be required to assess whether rotating stars would realistically be able to produce strong HeII emission lines (at the level of the recently detected Maiolino et al. 2024 Pop III galaxy candidate) late in the lifetimes of these stars, even if evolution were to follow the Yoon et al. (2012) models.
5 Impact on the 21 cm signal from cosmic dawn
The impact of Pop III stars on the cosmic 21-cm signal is multifold. During the epoch of cosmic dawn, the Lyman-band continuum flux of Pop III stars plays an important role through the Wouthuysen–Field (WF) effect, wherein Lyman- photons scatter off neutral hydrogen in the intergalactic medium and act to couple the 21-cm spin temperature to the kinetic temperature of the gas (Wouthuysen, 1952; Field, 1958). The scattering of Lyman-line photons also acts to raise the kinetic temperature of the intergalactic medium (e.g. Chuzhoy and Shapiro, 2007; Reis et al., 2021). At the same time, LW photons dissociate H2 molecules, suppressing further Pop III star formation in molecular cooling dark matter minihaloes (e.g. Fialkov et al., 2013; Visbal et al., 2014; Mirocha et al., 2018). During the epoch of reionization, Pop III stars that are still able to form ionize their surrounding gas, leading to a suppression in the 21-cm signal (e.g. Liu et al., 2025a)
In Fig. 6, we plot the lifetime-integrated Lyman-band photon production rates per ZAMS stellar baryon of our Pop III models in relation to the corresponding predictions presented by Gessey-Jones et al. (2022) for non-rotating Pop III stars, and those of Liu et al. (2025a) for rotating Pop III stars. Both of these are, like the Muspelheim ones, based on TLUSTY stellar atmosphere SEDs, but employ different assumptions concerning stellar evolution. The Gessey-Jones et al. (2022) and Liu et al. (2025a) models both trace evolution up to the end of main sequence under the assumption of no mass loss, whereas our models also include later evolutionary stages and – in the case of Yoon et al. (2012) models with rotation – substantial mass loss. Whereas the Gessey-Jones et al. (2022) models are based on non-rotating stars, the Liu et al. (2025a) models are based on Sibony et al. (2022) tracks for polytrope stars subject to chemically homogeneous evolution due to strong internal mixing. While Sibony et al. (2022) choose to remain agnostic about the physical reason for what would cause this mixing, fast rotation represents a strong candidate, as evident by the rotating Yoon et al. (2012) models, which in many cases also undergo chemically homogeneous evolution. For simplicity, we hereafter refer to the Liu et al. (2025a) models as rotating, in line with the discussion of these models by Liu et al. (2025a, b).
Differences in between the models at any given mass are within about a factor of two, with the rotating stars below 30 M⊙ consistently resulting in higher rates compared with their non-rotating counterparts due to their increased lifetimes. Interestingly for our discussion, we note a large difference compared to the Gessey-Jones et al. (2022) models at a mass of 20 M⊙, where the Yoon et al. (2012) models for rotating stars (which have the potential to boost the HeII1640 emission line equivalent width; Fig. 5) give rise to an that is higher by a factor of . Since it is important to establish whether differences of this amplitude could have any significant impact on 21-cm signatures from the pre-reionization Universe, we use the semi-numerical code 21cmSPACE (e.g. Fialkov et al., 2014; Reis et al., 2021; Gessey-Jones et al., 2022, 2025)222https://www.cosmicdawnlab.com/21cmSPACE to simulate the global 21-cm signal and the spherically averaged 21-cm power-spectrum at , under the extreme assumption of a single-star IMF where all Pop III stars have the same mass (a rough approximation of the case where the IMF is strongly peaked at this mass).
The simulations are performed on a cosmological box size of , and include effects such as baryon dark-matter streaming velocity (Fialkov et al., 2012), Lyman–Werner feedback (important during the cosmic dawn; Fialkov et al., 2013; Muñoz et al., 2022), Lyman- heating and scattering (Reis et al., 2021), photoheating feedback (Sobacchi and Mesinger, 2013; Cohen et al., 2016), X-ray heating (Fialkov et al., 2014; Fialkov and Barkana, 2014) and ionization from flexible spectra models (important during the epoch of reionization). In particular, for this work, we use the code version described in Liu et al. (2025a) where the effect of ionizing photons from Pop III stars was introduced. To put the Yoon et al. (2012) Pop III models into context, we also run 21cmSPACE with the Pop III models of Gessey-Jones et al. (2022) and Liu et al. (2025a).
We consider two models, with varying Pop III star formation efficiencies (i.e. fraction of gas converted in stars): Model 1 where Pop II SFR dominates at with and Model 2 where Pop III SFR is comparable to Pop II at with , as shown in Figure 8. In both models, we fix the Pop II star-formation efficiency to which generally matches observations at UV luminosity function observations at (Tacchella et al., 2018; Sipple and Lidz, 2024; Dhandha et al., 2025), and an star-formation averaging timescale of . We also fix Myr, which is the recovery time for Pop II star formation after the onset of Pop III supernovae (Magg et al., 2022a). To pinpoint the role of Lyman-band emissivities of Pop III stars in shaping the 21-cm signal, we switch off X-ray heating from Pop III stars. For Pop II, we adopt an X-ray luminosity tied to the star formation rate (SFR) of each halo, according to erg s-1 M yr-1 (Fragos et al., 2013b, a), which is consistent with constraints on the diffuse X-ray background, and HERA constraints on the 21-cm power spectrum at (Dhandha et al., 2025). Throughout, we assume instantaneous emission from Pop III stars, which is a reasonable assumption for stellar masses (Gessey-Jones et al., 2022).
Fig. 7 shows the resulting global 21-cm signal, and the power spectrum generated at a scale of cMpc-1, for six models: four models assuming a single-mass IMF of 20 M⊙ for the rotating and non-rotating Yoon et al. (2012) model, non-rotating Gessey-Jones et al. (2022) model, and a rotating Liu et al. (2025a) model, as well as two additional models of single-mass IMFs at 200 M⊙ for rotating and non-rotating Yoon et al. (2012) stars. The 200 M⊙ non-rotating Yoon et al. (2012) model is included here as an example of a type of star that, because of its high , could produce strong HeII1640 emission despite being non-rotating (Fig. 5).
In the case of the 21-cm power spectra, we also plot the expected 5 detection limit for 1080 hours of tracking-scan observations using the Square Kilometre Array (SKA-Low) with 6 hours of observation per night for 180 days. The derived detection limit considers both the thermal noise and the sample variance. To estimate the uncertainty in the power spectrum due to thermal noise, we used the package 21cmSense333https://github.com/rasg-affiliates/21cmSense/tree/main (Pober et al., 2013, 2014). We only considered the SKA-Low AA∗ array configuration (Sridhar et al., 2024) within a radius of 2 km from the array center and assumed the moderate foreground model in 21cmSense. The sample variance is calculated according to the equation
where N is the total number of counts and is the number of counts in the -bin. The total noise is then estimated as the sum of thermal noise and sample variance in quadrature:
5.1 The global 21-cm signal
The differences in the global signal between the Pop III stellar models can be explained by considering the epochs where the early onset of the global signal absorption trough is almost entirely dictated by the Lyman band emissivities (Fig. 6)444Although Lyman–Werner photons do contribute to suppression of star-formation in molecular cooling haloes, and we account for this., while the absorption trough at is also sensitive to the ionizing flux from the stars (Fig. 3).
During the cosmic dawn (), a higher Lyman-band flux leads to a stronger WF effect and more rapid onset of significant 21-cm absorption. This would be strongest for the rotating 20 M⊙ Yoon et al. (2012) model and weakest for the rotating 200 M⊙ Yoon et al. (2012) model, which indicates that the 21-cm signal observations can have the potential to distinguish the IMF of stellar populations (e.g. as also demonstrated in Gessey-Jones et al., 2025). The signal also shows differentiating features between the rotating and non-rotating 200 M⊙ Yoon et al. (2012) models, and a similar (but reversed) trend for the 20 M⊙ Yoon et al. (2012) models. In case of the rotating 20 M⊙ Liu et al. (2025a) model, since it exhibits a similar Lyman flux to the non-rotating 20 M⊙ Yoon et al. (2012), it shows similar behaviour at high redshifts. The difference in the 21-cm signal is of order few mK at the onset of the absorption signal up to () when Lyman-coupling is efficient (at for Model 1 and for Model 2).
During the epoch of heating and reionization (), the differences in these models is dominated by the ionization from Pop III stars (as explored in Liu et al., 2025a). The case of rotating models show convergence since they predict similarly strong ionizing fluxes across stellar masses (at M⊙), manifested as a suppressed absorption through in both Model 1 and Model 2. In the latter case, the differences are particularly large with a difference also seen in the redshift of the absorption signal as ionization becomes efficient. Indeed, this is reflected in the neutral hydrogen fraction shown in Fig. 9 with a reduction between the non-rotating and rotating cases (and thus, a proportional suppression of the signal). With the non-rotating models, the 20 M⊙ Gessey-Jones et al. (2022) model has the least efficient ionization, followed by 20 M⊙ Yoon et al. (2012) model (which shows similar absorption amplitude/timing), and lastly the 200. M⊙ Yoon et al. (2012) model.
5.2 The 21-cm power spectrum
The predicted high-redshift () 21-cm power spectra for Model 1 falls below the detection limit of the SKA, and close to the limit for Model 2. At lower redshifts (), where the power spectrum is deemed detectable according to our error model, the differences between stellar models are marginal in Model 1 (especially if considering degeneracies with other astrophysical parameters that we keep fixed here). However, there are more pronounced and detectable differences in Model 2 which come from the ionization signature of the rotating Pop III stars, as in the global signal. In these cases, as the Universe is rapidly ionized, the ionized bubbles imprint boosted fluctuations in the power spectrum at . The non-rotating cases show a similar trend to the global signals, where the 20 M⊙ Gessey-Jones et al. (2022) model shows the smallest ionization peak of mK2, while the 20 M⊙ Yoon et al. (2012) model and 200 M⊙ Yoon et al. (2012) model show higher peaks by and .
Overall, we conclude that the dramatic difference seen between the He+-ionizing rates (Fig. 4) and HeII1640 equivalent widths (Fig. 5) can imprint some detectable 21-cm signatures. The differences are modest for low Pop III star-formation efficiency (Model 1), potentially only detectable with high-accuracy global signal experiments (noise levels of mK) at . For high Pop III star-formation efficiency (Model 2, which shows early reionization), the differences imprinted in the 21-cm power spectrum at are more dramatic and detectable by SKA.
6 Discussion
6.1 Blackbody versus stellar atmosphere SEDs
When comparing our results with the models based on blackbody spectra presented by Klessen and Glover (2023) in Fig. 3, we find a large discrepancy is in the He+ ionizing rates for stars , where blackbody spectra result in rates orders of magnitude higher than those based on stellar atmosphere spectra. These differences, also echoed in our Murphy et al. (2021a) comparison of momentary for stellar atmosphere spectra versus blackbody spectra in Fig. 2, are due to the different shapes of blackbody and stellar atmosphere SEDs and the lack of continuum breaks in the former case, as shown in Fig. 1. Hence, blackbody spectra should not be used to estimate Pop III photon production rates or fluxes in the He+ ionizing energy range.
We also note significant deviations in the other UV bands for stars . While Murphy et al. (2021b) did comment on the reliability of the blackbody approximation for the Murphy et al. (2021b) tracks (through a comparison to the results of Schaerer 2002), this offset was not noted since the Murphy et al. (2021b) study of UV photon production rates considered only stars of masses from the Murphy et al. (2021a) set. The blackbody approximation from their lower-mass tracks is adopted by Klessen and Glover (2023) which overestimates the ionizing rates for all low mass stars and slightly underestimates the LW rates for these same stars. However, as long as the Pop III IMF has a characteristic mass – which is often argued to be the case (e.g. Hartwig et al., 2024; Liu et al., 2024) – this discrepancy should not be a major source of error, as the ionizing photon production in that case would be dominated by the more massive stars.
6.2 The Lyman–Werner emissivities of Pop III stars
The LW photon production rates of Pop III stars are important for establishing the LW radiation background during cosmic dawn and for the quenching of subsequent star formation in low-mass haloes. When the lifetime-integrated is expressed as LW photons per stellar baryon, there are two distinct values that are frequently featured in the literature: 4800 and . Our models do not favour either of these, but instead favour values in the intermediate range to (see Fig. 3; values in this figure can be converted to log LW photons per stellar baryon by subtracting ). The lower literature value (4800) stems from Barkana and Loeb (2005), based on Pop III spectra from Bromm et al. (2001) and an adopted power-law slope of the spectrum, whereas the higher value () is usually attributed to Schaerer (2002). Indeed, when the Schaerer (2002) Table 4 values are converted into these units, they peak at LW photons per baryon at a Pop III stellar mass of 60 M⊙. However, as shown in Fig. 3, the Schaerer (2002) are higher than ours at by up to a factor (0.7 dex). Mas-Ribas et al. (2016) also note difficulties in reproducing the Schaerer (2002) LW rates and suggest that the Schaerer (2002) LW rates may have been integrated from 11.2 eV (wavelength 1107 Å) to infinity, rather than to 13.6 eV (wavelength 912 Å), which would lead to boosted rates. Lifetime-integrated LW rates per stellar baryons in the range for Pop III stars are also reported in the recent paper by Costa et al. (2025), and in that case the LW rates were indeed integrated from 11.2 eV to infinity (Guglielmo Costa, private communication). After correcting the Costa et al. (2025) limit from 11.2 to 13.6 eV, their values agree with ours. Our values are moreover in good agreement with those recently presented by Liu et al. (2025a) for the non-rotating Pop III tracks of Gessey-Jones et al. (2022) and the rotating Pop III tracks of Sibony et al. (2022). Hence, our recommendation is to use values for the lifetime-integrated LW photon counts per stellar baryons of to .
6.3 Viability of the stellar atmosphere models
The results in this paper are based on SED models for Pop III stars, which at K make use of TLUSTY (Hubeny and Lanz, 1995) stellar atmosphere models for plane-parallel, hydrostatic stars with primordial surface composition. While certainly more realistic than blackbody SEDs, such stellar atmosphere models may still represent an oversimplification that could impact the emitted ionizing photon production rates of Pop III stars. The surface He/H composition of Pop III stars may change over the course of their lifetimes, and the surface may also get polluted by metals produced within the star, due to either chemically homogeneous evolution for a rapidly rotating star (e.g. Yoon et al., 2012; Liu et al., 2025a), but also due to dredge-ups for a non-rotating star (e.g. Volpato et al., 2023).
While Schaerer (2002) notes that the impact of the H/He level on the stellar atmosphere spectra is small, this statement likely does not apply to models for rotating stars with helium mass fractions close to unity (Sibony et al., 2022). Liu et al. (2025a) do take the evolving helium abundance into account in their modelling of Pop III stars, and while they note that this helps the stellar atmosphere model converge, no analysis of the impact of the pure-helium atmosphere on the emerging stellar atmosphere spectra was presented in their paper.
Indeed, as shown in Zackrisson et al. (2024), the set of TLUSTY stellar atmosphere models used by Muspelheim (with fixed H/He fractions) are – due to convergence problems – sometimes unable to reach the low surface gravities of all Pop III stellar evolutionary tracks. This is for instance the case of the K state of the rotating 20 M⊙ Yoon et al. (2012) stars highlighted in Section 4 in relation to its potential to produce a very strong HeII1640 emission line. These 20 M⊙ rotating stars lie below the last converged grid data point. However, based on how the He+-flux changes with within the converged grid at these , this is likely to have a very minor effect on .
Rotational mixing and surface composition can also be affected by the presence of a companion Pop III star in a binary system (Song et al., 2020). Rapid rotation could furthermore drive a mechanical wind, and metals at the surface of Pop III stars may also trigger line-driven winds (e.g. Liu et al., 2021; Jeena et al., 2023). The effects of winds on the emerging ionizing photon production rates of Pop III stars were explored by Schaerer (2002), who concluded that the ionizing photon rates would not be substantially affected compared to TLUSTY stellar atmosphere spectra in their models, even in their high mass-loss case. However, Schaerer (2002) also notes that for sufficiently dense winds, the He+ flux could potentially get substantially suppressed (as demonstrated by Chen et al. 2015 for metal-enriched stars). It may therefore be interesting to revisit this issue in the future, using stellar atmosphere spectra with updated wind models for Pop III stars. However, it should be noted that if winds are launched only late in the evolution of Pop III stars, this is unlikely to matter for the time-integrated emission of ionizing photons into the surrounding medium, unless the Pop III stars also exhibit high in such stages.
Finally, rapidly rotating stars would not be able to retain spherical shapes, and would exhibit different and surface gravities at their poles compared to their equators, thereby exhibiting anisotropic SEDs. While orientations would be randomized in a large population of such stars, mechanical mass-loss due to rotation may also preferentially happen in the direction perpendicular to the rotation axis, which could introduce a systematic bias in the UV emissivities seen by the surrounding interstellar and intergalactic medium, compared with the simplified models used here.
6.4 Binary Pop III stars
In this paper, we have exclusively considered models for single stars, even though Pop III stars in simulations often form binary and higher-order systems (e.g. Stacy and Bromm, 2013; Sugimura et al., 2020). As shown by Tsai et al. (2023), binary interactions can boost the UV fluxes of Pop III stars (in the case of lifetime-integrated fluxes by up to a factor of in the LW band, which is the strongest effect seen in their models). However, binary interactions are also expected to affect stellar rotation rates (e.g. de Mink et al., 2013), which by themselves affect stellar evolution and consequently these fluxes. Given the many open questions regarding the stellar initial mass function, initial rotation rate and binary interactions of Pop III stars, models of the UV flux from these stars are likely to suffer from significant uncertainties until the arrival of better observational constraints. As rotation and binary interactions affect the end states of Pop III stars, such constraints may for instance come from the study of Pop III supernova yields in second generation stars or gas absorption systems, from direct observations of Pop III supernovae or gamma-ray bursts (and their afterglows), or from future gravitational wave detectors.
7 Conclusions
In this work, we have presented a set of publicly available models for the UV photon production rates of single Pop III stars, with and without rotation, across a wide range of stellar masses. We have investigated how the photon production rates are affected by the adoption of blackbody spectra compared with stellar atmosphere spectra. We have also investigated and compared the resulting lifetime-integrated emissivities of several different sets of Pop III tracks. Lastly, for selected models, we have explored the potential role of stellar rotation on the equivalent width of the HeII1640 emission line, on the 21-cm global signal, and on the the spherically averaged 21-cm power spectrum.
Our conclusions can be summarized as follows:
1.
The assumption of blackbody spectra generally overestimate He+ ionizing rates, in some cases by more than an order of magnitude, depending on the of the star.
2.
Based on the stellar evolutionary tracks of Yoon et al. (2012), which take rotation and magnetic fields into account and predict evolution to K at the end stages of evolution, we show that rotation significantly increases the total He+ ionizing photon production of Pop III stars in the range 20 to 100 M⊙ compared with stars without rotation.
3.
If gas can be retained with high-redshift Pop III systems dominated by rotating stars, then such systems may be able to produce very strong HeII1640 emission with rest-frame equivalent widths comparable to or higher than systems dominated by Pop III stars.
4.
Differences between rotating and non-rotating Pop III stars in 21-cm observables like the 21-cm global signal and the 21-cm power spectrum are modest for sub-per cent level Pop III star-formation efficiency, and more pronounced for per cent Pop III star-formation efficiency, with potential for detectability with the SKA. The differences are of order per cent for the global signal (at ) and a factor of for the ionization peak in the power spectrum (at ) due to the extreme ionizing flux of rotating stars.
5.
We highlight that two disparate values for the Lyman–Werner rates of Pop III stars have been used in the previous literature on 21-cm cosmology and high-redshift star formation ( vs LW photons per stellar baryon). Our models instead favour a value around LW photons per stellar baryon, in line with other recent estimates (e.g. Liu et al., 2025a).
Acknowledgements
This manuscript is based on work initially conducted as part of the first author’s Master’s thesis Wasserman (2025). The work has been expanded upon and revised in collaboration with the co-authors for this publication. EZ acknowledges project grant 2022-03804 from the Swedish Research Council (Vetenskapsrådet). This work was supported in part by the International Centre for Theoretical Sciences (ICTS) for participating in the program - Radio Cosmology and Continuum Observations in the SKA Era: A Synergic View. (code: ICTS/radiocoscon2025/04). An anonymous referee provided useful comments that helped improve the quality of the manuscript.
Data Availability
Publicly available grids of Muspelheim models (SEDs, ) and both lifetime-integrated and lifetime-average photon production rates are available from: https://www.astro.uu.se/˜ez/muspelheim/muspelheim.html
Appendix A Extra tables and figures
Table 1 shows the average lifetime photon production rates for the Yoon et al. (2012) and Murphy et al. (2021a) models used in this work. Fig. 9 shows the neutral fractions from the 21cmSPACE simulations in Fig. 7.
| Source | Rotation | Mass | Lifetime | |||||
| Yoon+12 | 0.0 | 10 | 47.71 | 47.23 | 41.98 | 47.57 | 47.75 | 2.61E+07 |
| Yoon+12 | 0.0 | 20 | 48.66 | 48.31 | 43.79 | 48.20 | 48.37 | 1.02E+07 |
| Yoon+12 | 0.0 | 30 | 49.10 | 48.77 | 44.87 | 48.54 | 48.70 | 6.81E+06 |
| Yoon+12 | 0.0 | 60 | 49.68 | 49.39 | 47.66 | 49.00 | 49.17 | 4.16E+06 |
| Yoon+12 | 0.0 | 100 | 50.02 | 49.74 | 48.54 | 49.29 | 49.46 | 3.27E+06 |
| Yoon+12 | 0.0 | 150 | 50.27 | 50.00 | 48.98 | 49.54 | 49.71 | 2.70E+06 |
| Yoon+12 | 0.0 | 200 | 50.40 | 50.14 | 49.19 | 49.66 | 49.83 | 2.64E+06 |
| Yoon+12 | 0.0 | 300 | 50.59 | 50.33 | 49.45 | 49.85 | 50.01 | 2.40E+06 |
| Yoon+12 | 0.0 | 500 | 50.83 | 50.57 | 49.73 | 50.08 | 50.25 | 2.18E+06 |
| Yoon+12 | 0.4 | 10 | 47.56 | 46.98 | 41.53 | 47.57 | 47.75 | 2.74E+07 |
| Yoon+12 | 0.4 | 20 | 48.88 | 48.63 | 47.48 | 48.12 | 48.28 | 1.52E+07 |
| Yoon+12 | 0.4 | 30 | 49.26 | 49.03 | 48.00 | 48.40 | 48.56 | 9.51E+06 |
| Yoon+12 | 0.4 | 60 | 49.79 | 49.59 | 48.75 | 48.82 | 48.98 | 5.28E+06 |
| Yoon+12 | 0.4 | 100 | 50.12 | 49.94 | 49.21 | 49.10 | 49.26 | 3.93E+06 |
| Yoon+12 | 0.4 | 150 | 50.36 | 50.18 | 49.48 | 49.32 | 49.47 | 3.28E+06 |
| Yoon+12 | 0.4 | 200 | 50.52 | 50.32 | 49.59 | 49.54 | 49.69 | 2.97E+06 |
| Yoon+12 | 0.4 | 300 | 50.69 | 50.48 | 49.68 | 49.74 | 49.90 | 2.55E+06 |
| Yoon+12 | 0.4 | 500 | 50.87 | 50.62 | 49.74 | 50.07 | 50.25 | 2.23E+06 |
| Murphy+21 | 0.0 | 9 | 47.68 | 47.13 | 43.98 | 47.30 | 47.61 | 2.00E+07 |
| Murphy+21 | 0.0 | 12 | 48.10 | 47.63 | 44.86 | 47.48 | 47.83 | 1.93E+07 |
| Murphy+21 | 0.0 | 15 | 48.23 | 47.90 | 45.47 | 47.55 | 48.01 | 1.40E+07 |
| Murphy+21 | 0.0 | 20 | 48.84 | 48.40 | 46.19 | 48.07 | 48.26 | 1.03E+07 |
| Murphy+21 | 0.0 | 30 | 49.07 | 48.82 | 46.95 | 47.91 | 48.58 | 6.64E+06 |
| Murphy+21 | 0.0 | 40 | 49.45 | 49.16 | 47.40 | 47.87 | 48.81 | 5.21E+06 |
| Murphy+21 | 0.0 | 60 | 49.81 | 49.52 | 47.94 | 48.34 | 49.09 | 4.03E+06 |
| Murphy+21 | 0.0 | 85 | 50.15 | 49.79 | 48.37 | 49.31 | 49.34 | 3.40E+06 |
| Murphy+21 | 0.0 | 120 | 50.38 | 50.02 | 48.72 | 49.60 | 49.54 | 3.05E+06 |
| Murphy+21 | 0.4 | 9 | 47.57 | 47.06 | 41.73 | 47.47 | 47.65 | 2.38E+07 |
| Murphy+21 | 0.4 | 12 | 47.93 | 47.52 | 42.42 | 47.74 | 47.92 | 2.25E+07 |
| Murphy+21 | 0.4 | 15 | 48.26 | 47.89 | 43.02 | 47.94 | 48.11 | 1.65E+07 |
| Murphy+21 | 0.4 | 20 | 48.64 | 48.30 | 43.73 | 48.18 | 48.34 | 1.19E+07 |
| Murphy+21 | 0.4 | 30 | 49.09 | 48.77 | 44.77 | 48.49 | 48.65 | 7.60E+06 |
| Murphy+21 | 0.4 | 40 | 49.36 | 49.05 | 45.92 | 48.70 | 48.86 | 5.76E+06 |
| Murphy+21 | 0.4 | 60 | 49.70 | 49.39 | 47.50 | 49.01 | 49.18 | 4.41E+06 |
| Murphy+21 | 0.4 | 85 | 49.94 | 49.64 | 48.20 | 49.24 | 49.41 | 3.60E+06 |
| Murphy+21 | 0.4 | 120 | 50.17 | 49.89 | 48.65 | 49.43 | 49.59 | 3.14E+06 |